A Researcher would like to test the effectiveness of a newly developed growth hormone. The Researcher knows that under normal circumstances laboratory rats reach an average weight of μ = 950 Grams at 10 Weeks of age. The Distribution of weights is normal with σ = 30. A random sample of n = 25 Newborn rats is obtained, and the hormone is given to each rat. When The rats in the sample reach 10 Weeks old, each rat is weighed. The Mean weight for this sample is M = 974.
a. Identify the independent and the dependent variables for this study.
b. Assuming a two-­‐tailed test, state the null hypothesis in a sentence that includes the independent variable and the dependent variable.
c. Using symbols, state the hypotheses (H0 And H1) For the two tailed test.
d. Sketch the appropriate distributions, and locate the critical region for α =.05.
e. Calculate the test statistic (z-­‐score) For the sample.
f. What decision should be made about the null hypothesis, and what decision should be made about the effect of the hormone?

Answer by stanbon(75887)

A Researcher would like to test the effectiveness of a newly developed growth hormone.
The Researcher knows that under normal circumstances laboratory rats reach an average weight of μ = 950 Grams at 10 Weeks of age.
The Distribution of weights is normal with σ = 30.
A random sample of n = 25 Newborn rats is obtained, and the hormone is given to each rat.
When The rats in the sample reach 10 Weeks old, each rat is weighed. The Mean weight for this sample is M = 974.
------------------------------------------
a. Identify the independent and the dependent variables for this study.
Independent: amt of hormone
dependent: weight of the rats
------------------------------------------
b. Assuming a two-­‐tailed test, state the null hypothesis in a sentence that includes the independent variable and the dependent variable.
Hormones do or do not make a difference in rat weights.
-------------------------------
c. Using symbols, state the hypotheses (H0 And H1) For the two tailed test.
Ho: u(no hor)-u(hor) = 0
Ha: u(no hor)-u(hor) is not equal to 0
-------------------------------
d. Sketch the appropriate distributions, and locate the critical region for α =.05.
+-invT(0.975 with df=24) = +-2.0639
-------------------------------
e. Calculate the test statistic (z-­score) For the sample.
t(950-974)= -24/(standard deviation)
Note: No standard deviation for the sample was posted.
-------------------------------
f. What decision should be made about the null hypothesis, and what decision should be made about the effect of the hormone?
Note: No decision because test statistic cannot be calculated.
=====================================
Cheers,
Stan H.